OpenStudy (anonymous):

The sum of two numbers is 60 and their difference is 12 find the two numbers

OpenStudy (anonymous):

Hey Jt, you know sum means addition so let two numbers equal 60 such as A+B=60. Let their difference (or subtraction) be 12 so A-B=12. Solve for A and B then you'll get A=36 and B=24. Hope this helps!

OpenStudy (marco26):

x+Y = 60 X-Y=12 X=Y+12 Y+12+Y=60 2Y=60-12 Y=24 X=24+12 X=36